# Forcing Diffuse Distributions out of Language Models

### v1

@@ Abstract
Despite being trained specifically to follow user instructions, today’s language models perform poorly when instructed to produce random outputs. For example, when prompted to pick a number uniformly between one and ten Llama-2-13B-chat disproportionately favors the number five, and when tasked with picking a first name at random, Mistral-7B-Instruct chooses Avery 40 times more often than we would expect based on the U.S. population. When these language models are used for real-world tasks where diversity of outputs is crucial, such as language model assisted dataset construction, their inability to produce diffuse distributions over valid choices is a major hurdle. In this work, we propose a fine-tuning method that encourages language models to output distributions that are diffuse over valid outcomes. The methods we introduce generalize across a variety of tasks and distributions and make large language models practical for synthetic dataset generation with little human intervention. [^1] Code and data are available at https://github.com/y0mingzhang/diffuse-probabilities.
@@ 1 Introduction
Figure 1: Language models do not produce diffuse probabilities. The output distributions of baseline Gemma, Llama-2, and Mistral models deviate from what we expect from natural/random distributions. Our tuning method addresses this issue by diffusing the output distribution over valid candidates. In each plot, the horizontal axis is sorted in descending order by probability of the specific output.
@@
Language models are extremely bad at producing random outputs when users want them to. Even when prompts are carefully constructed with instructions that encourage randomness, both state-of-the-art open-source and industry language models output very low-entropy distributions over the valid options. Beyond Dungeons & Dragons, there are many practical applications where diversity across valid options is crucial for language model outputs. For example, when language models are used to answer multiple choice or Likert-scale questions, a priori each option should be equally likely. When they are used for synthetic dataset construction, such as for synthetic biographies (Maini et al., 2024; Yuan et al., 2021) or instruction-tuning train sets (Wang et al., 2023), diversity in the generations is crucial but arduous to achieve through mere prompt hacking.
In this work, we examine just how off language model generations are from user expectations of randomness and diversity. We then show how language models can be fine-tuned to produce diffuse distributions over valid options, without sacrificing generation quality. Our method supports tasks where the sample set of valid options is not easily enumerated, and we show that models fine-tuned to produce diffuse probabilities for one set of tasks generalize to other, very different tasks. This generalization allows us to promote diversity in complex settings, such as synthetic dataset generation. On the task of generating a dataset of synthetic biographies (similar to the dataset created by Maini et al. (2024) for benchmarking language model unlearning), our method generates four times as many unique first names, three times as many unique birth places, and 1.5 times as many unique careers as the baseline model, all without any need for complex prompt engineering, decoding strategy tweaking or manual re-writing.
@@ 2.1 Problem setting
In generation, we are mostly interested in computing the probability of a multi-token target $\mathbf{y}=[y_{1},y_{2},\dots,y_{n}]\in\mathcal{V}^{n}$ (e.g., $\mathbf{y}$ could be a two-digit number or the biography of a person). An autoregressive language model factors the probability of a sequence $\mathbf{y}$ into the product of probabilities of individual tokens:222$\oplus$ indicates concatenation.
@@ 2.2 Measuring output distributions
How do we measure whether a language model produces diffuse probabilities? We consider two methods, one for cases where the sample space is enumerable, and another for when the full sample space is unknown.
@@ 3.1 Distribution matching as language model fine-tuning
Since the model distribution $\mathbf{p}_{\theta}$ is defined on sequences, this optimization relies on a mild condition that $\mathcal{T}$ is prefix-free, namely no target sequence $\mathbf{t}$ should be the prefix of another target sequence $\mathbf{t}^{\prime}$. This condition is necessary since we would otherwise always have $\mathbf{p}_{\theta}(\mathbf{t}^{\prime}\,|\,\mathbf{x})<\mathbf{p}_{\theta}(\mathbf{t}\,|\,\mathbf{x})$, and it would be impossible to, say, make a model output identical probability for $\mathbf{t}$ and $\mathbf{t}^{\prime}$. [^3] We can simply append an end-of-sequence token to each target if $\mathcal{T}$ is not already prefix-free.
Notice that minimizing the objective in Equation 5 is equivalent to minimizing the KL-divergence between the model and the ground truth distribution $\mathbf{p}^{\star}$:
@@ 3.2 Parameter-efficient Fine-tuning
We use LoRA (Hu et al., 2021) to fine-tune models in a parameter-efficient way. This technique relies on optimizing low-rank additions to weight matrices, making fine-tuning large language models (with $\geq 7$B parameters) practical. It has the added benefit that LoRA tuning tends to perform almost as well as full model fine-tuning and generalizes beyond training instances (Malladi et al., 2023). From an implementational standpoint, LoRA requires instantiating a new pair of relatively small matrices for every weight matrix in the network. Consider one fully connected layer parameterized by $W\in\mathbb{R}^{d\times d}$. We instantiate $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times d}$ where $r\ll d$ and reparametrize the layer as $W+BA$. When fine-tuning the model, we exclusively optimize the weights of $A$ and $B$ rather than taking gradient steps on $W$, enabling us to fine-tune large models such as Llama-2-13B with only a small fraction of the memory overhead. [^4] Empirically, we find that $r=4$ is sufficiently expressive, which we use across all experiments. On our models, this corresponds to tuning less than 0.1% of all model parameters.
@@ 4 Results on Simple Tasks
We begin by showing how state-of-the-art language models fail to produce diffuse distributions on simple tasks, even when a user’s prompt requests randomness. We consider two tasks, Baby Names and Random Number Generation (RNG). In Baby Names, the model is instructed to generate a random baby name in English. In RNG, we ask the model to generate a random number between 1 and 10. The tasks have the following prompts:
@@
We consider three state-of-the-art language models: Gemma-7B-instruct (Google, 2024), Llama-2-13B-chat (Touvron et al., 2023) and Mistral-7B-instruct (Jiang et al., 2023). We use chat/instruct versions of the models, which are fine-tuned on dialog data to make them more suitable than base models for our test cases, which require instruction following. In all experiments, we sample from models using a temperature of one, which is equivalent to sampling from the maximally diffuse, unmodified model distributions. [^6] We could technically set temperature to beyond one to smooth out the model distribution further, but this is rarely done in practice due to degeneration. We also note that the baseline models need to be sampled at a temperature $>10$ to match the entropy of our fine-tuned models in Section 4.2. For each task, we sample 1,000 generations from the model and report coverage and entropy statistics of the empirical distributions of model generations.
@@ 4.1 Language models fail to follow instructions for randomness
Figure 1 shows output distributions of several untuned models for Baby Names and RNG. Despite the prompts asking for randomness, these output distributions are extremely imbalanced.
For RNG, it is easy to see that the empirical distribution is far from the uniform distribution we would expect. For Baby Names, it is less clear what we should expect from the model. Do we want the generated names to be a uniform sample from the set of all names? Or do we want the names to reflect the distribution of baby names among recently born babies in the Unites States? To address this, we introduce the concept of an natural distribution for each task. For RNG, this is simply the uniform distribution; for Baby Names, we compare to a sample we draw from data on names used for babies born in the United States in the year 2022.777Using birth data from the US Social Security Administration. We note, that we are only using the natural distributions as reasonable baselines for comparison.
@@ 4.2 Optimizing for Baby Names and RNG diversifies generation
In these experiments, we see our methods can very effectively align the model distribution with the ideal distributions. With RNG, tuned models produce near-uniform distributions over the valid choices with full coverage and near perfect entropy. [^8] Coverage and entropy results are reported in Appendix B.1. With Baby Names, where the training set covers only a small portion of the possible outputs, we see an encouraging indication of generalization: while the coverage remains smaller than that of the natural distribution, it is over twice the baseline for all models and the majority of outputs are not among targets in the fine-tuning set.
@@
While RNG and Baby Names serve as toy problems, these settings have a major limitation—the fine-tuned models are evaluated on the same prompts and sample spaces in training. To further test the generalization performance of our models, we also consider settings where the prompt and sample space differ between training and evaluation. We fine-tune Llama-2 to produce diffuse probabilities over two ranges of random integers—one to ten and one to 100—and test its generalization to different instruction formats and unseen sample spaces (see Appendix A.1). In Figure 2(a), we observe encouraging generalization trends: the tuned model produces near uniform distributions for unseen prompt formats over number ranges not in the fine-tuning set, for example from 1 to 45. When we vary the size of the random sample space of RNG (e.g., 154 to 204), the tuned model still produces substantially higher entropy distributions than the baseline model (Figure 2(b)). With these first experiments, we demonstrate that tuning models for diffuse probabilities is a promising method for increasing generation diversity.
@@ 5 Generalization Across Tasks
We use leave-one-out experiments on a set of six tasks to show that models tuned for diffuse probabilities do, in fact, have strong transferrability to tasks unseen during tuning. We consider the following six tasks. Each task has several associated prompts, one of which is shown below:
@@
In Figure 3, we report in-distribution (ID) and out-of-distribution (OOD) results, which correspond to tuning sets that include or exclude the particular task, respectively. The results show a convincing trend: for all three models, our tuning method led to substantial improvements in entropy over the baselines, even when the task was held out from the tuning set. [^9] Coverage results show similar trends, and we report them in Table 5, Appendix B.2. Another interesting observation is that the baseline Mistral model consistently produces more diffuse distributions than the baseline Gemma and Llama-2 models, but after tuning, all three models have comparable entropy.
In two of the tasks (Countries and Fruits), we observe sizable generalization gaps between in-distribution and out-of-distribution entropy, which suggest that task-specific tuning remains useful, especially when we can come up with a reasonbly diverse set of generation targets for fine-tuning. However, coming up with a large enough target set isn’t always easy. In these cases, we rely on the generalization of the models trained on a diverse set of tasks. For example, in Occupations, the authors could only come up with a small set of 17 professions that start with the letter “A,” and Llama-2 and Mistral (not trained on Occupations) generalize out-of-distribution to occupations beyond what we provide in the fine-tuning set. [^10] For example, Llama-2 generated “Aromatherapist”, and Mistral generated “Agronomist.” Neither is among the fine-tuning targets of Occupations. Notably, our fine-tuning method does not substantially change the general capabilities (e.g., writing and reasoning) of the models demonstrated by evaluations on MT-Bench (see Appendix B.5), making our method compatible with tuning general-purpose large language models.
@@ 6 Constructing More Diverse Synthetic Datasets
We first consider both a Llama-2-13B model fine-tuned on all tasks in Section 3 and not specifically for biographies (Tuned-OOD), and compare its generations against the baseline Llama-2 model over 1000 samples. In Figure 4, we report coverage results for categorical attributes (e.g., Birth year), and normalized unigram diversity for achievements and over the entire biography. [^11] We report prompt used for biography generation and side-by-side qualitative examples in Appendix A In Table 1, we report the most frequently generated values for categorical attributes, along with the frequencies out of 1000 generations. [^12] A table of coverage and entropy statistics of generated biographies can be found in Appendix B.3.
The results on the baseline Llama-2 model indicate that the biases we observe in Section 4.1 towards certain names and numbers expectedly show up in generated data. For example, out of 1000 biographies generated by the baseline Llama-2 model, 284 have the first name ’Evelyn,’ and 966 are females. Such a high level of repetition in the generation makes the resulting dataset basically unusable for any downstream task without substantial human intervention. In constrast, our model significantly improves the diversity of generated biographies, despite not trained specifically for generating biographies. For example, we see an $>$2X increase in coverage for most categorical attributes. There are also substantial improvements in generation diversity for achievements and over the entire biography (Figure 4(b)), although our training did not optimize for open-ended text generation. The top five most frequent values for attributes (Table 1) help contextualize this improvements in diversity: there are significant reductions in biases towards certain attribute values. E.g., the frequency of the name “Evelyn” decreased by 25X, and the birth year 1985 by over 3X.
6.1 Controlling distributions of categorical attributes
(a) Coverage results for categorical attributes.
(b) Unigram diversity of biographies.
At a comparable level of generation diversity to the Tuned-OOD model (Figure 4), the Tuned-ID model is able to generate biographies with much more balanced distributions of categorical attributes: Table 1 shows that both gender are birth years are roughly uniformly distributed, and a wider range of birth places are produced by the model. [^13] See maps of generated birth places for all three models in Figure 5, Appendix B. Crucially, the model remains highly diverse on open-ended generation of achievements even when being trained exclusively on categorical attributes. This result highlights the potential in extending our fine-tuning method towards improving diversity in open-ended text generation.

### v2

@@ Abstract
Despite being trained specifically to follow user instructions, today’s instruction-tuned language models perform poorly when instructed to produce random outputs. For example, when prompted to pick a number uniformly between one and ten Llama-2-13B-chat disproportionately favors the number five, and when tasked with picking a first name at random, Mistral-7B-Instruct chooses Avery 40 times more often than we would expect based on the U.S. population. When these language models are used for real-world tasks where diversity of outputs is crucial, such as language model assisted dataset construction, their inability to produce diffuse distributions over valid choices is a major hurdle. In this work, we propose a fine-tuning method that encourages language models to output distributions that are diffuse over valid outcomes. The methods we introduce generalize across a variety of tasks and distributions and make large language models practical for synthetic dataset generation with little human intervention. [^1] Code and data are available at https://github.com/y0mingzhang/diffuse-distributions.
@@ 1 Introduction
Figure 1: Language models do not produce diffuse probabilities. The output distributions of untuned Gemma, Llama-2, and Mistral deviate from what we would expect from natural/random distributions. Our tuning method addresses this by diffusing the output distribution over valid candidates. The horizontal axes above are sorted in descending order by probability of the output.
@@
Instruction-tuned language models are extremely bad at producing random outputs when users want them to. Even when prompts are carefully constructed with instructions that encourage randomness, both state-of-the-art open-source and industry language models output very low-entropy distributions over the valid options. Beyond Dungeons & Dragons, there are many practical applications where diversity across valid options is crucial for language model outputs. For example, when language models are used to answer multiple choice or Likert-scale questions, a priori each option should be equally likely. When language models are used for synthetic dataset construction, such as for synthetic biographies (Maini et al., 2024; Yuan et al., 2021) or instruction-tuning training sets (Wang et al., 2023), diversity in the generations is essential but arduous to achieve through mere prompt hacking.
In this work, we examine how far off language model generations are from user expectations of randomness and diversity. While our experiments focus on instruction-tuned models such as Llama-2-chat, our findings do generalize to pre-trained language models in general. [^2] For example, when GPT-2 (Radford et al., 2019) is prompted with “She rolled the 20-sided die. It landed on the number”, the output distribution is similarly far from uniform. We then show how language models can be fine-tuned to produce diffuse distributions over valid options, without sacrificing generation quality. Our method supports tasks where the sample set of valid options is not easily enumerated, and we show that models fine-tuned to produce diffuse probabilities for one set of tasks generalize to other, very different tasks without degrading model utility.
This generalization allows us to promote diversity in complex settings, such as synthetic dataset generation. On the task of generating a dataset of synthetic biographies (similar to the dataset created by Maini et al. (2024) for benchmarking language model unlearning), our method generates four times as many unique first names, three times as many unique birth places, and 1.5 times as many unique careers as the untuned baseline model, all without any need for complex prompt engineering, decoding strategy tweaking or manual re-writing.
@@ 2.1 Problem setting
In generation, we are interested in computing the probability of a potentially multi-token target $\mathbf{y}=[y_{1},y_{2},\dots,y_{n}]\in\mathcal{V}^{n}$. For example, $\mathbf{y}$ could be a two-digit number or the biography of a person. A language model factors the probability of a sequence $\mathbf{y}$ into the product of probabilities of individual tokens ($\oplus$ here indicates concatenation):
@@ 2.2 Measuring output distributions
How do we measure if an LM produces diffuse probabilities? We consider one method for cases where the sample space is enumerable, and another for when the full sample space is unknown.
@@ 3.1 Distribution matching as language model fine-tuning
This objective is, in fact, the cross-entropy loss ubiquitous in language modeling (Berger et al., 1996) extended to multiple target sequences (Edunov et al., 2018). Since the model distribution $\mathbf{p}_{\theta}$ is defined on sequences, this optimization relies on a mild condition that $\mathcal{T}$ is prefix-free, namely no target sequence $\mathbf{t}$ should be the prefix of another target sequence $\mathbf{t}^{\prime}$. This condition is necessary since we would otherwise always have $\mathbf{p}_{\theta}(\mathbf{t}^{\prime}\,|\,\mathbf{x})<\mathbf{p}_{\theta}(\mathbf{t}\,|\,\mathbf{x})$, and it would be impossible to, say, make a model output identical probability for $\mathbf{t}$ and $\mathbf{t}^{\prime}$. [^3] We can simply append an end-of-sequence token to each target if $\mathcal{T}$ is not already prefix-free. Notice that minimizing the objective in Eq. 5 is equivalent to minimizing the KL-divergence between the model and the ground truth distribution $\mathbf{p}^{\star}$:
@@ 3.2 Parameter-efficient Fine-tuning
We use LoRA (Hu et al., 2021) to fine-tune models in a parameter-efficient way. This technique relies on optimizing low-rank additions to weight matrices, making fine-tuning large language models (with $\geq 7$B parameters) practical. It has the added benefit that LoRA tuning tends to perform almost as well as full model fine-tuning and generalizes beyond training instances (Malladi et al., 2023). From an implementational standpoint, LoRA requires instantiating a new pair of relatively small matrices for every weight matrix in the network. Consider one fully connected layer parameterized by $W\in\mathbb{R}^{d\times d}$. We instantiate $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times d}$ where $r\ll d$ and reparametrize the layer as $W+BA$. When fine-tuning the model, we exclusively optimize the weights of $A$ and $B$ rather than taking gradient steps on $W$, enabling us to fine-tune large models such as Llama-2-13B with only a small fraction of the memory overhead. [^4] Empirically, we find that $r=4$ is sufficiently expressive, which we use across all experiments. On our models, this corresponds to tuning less than 0.1% of all model parameters. We report training details in Appendix A.
@@ 4 Results on Simple Tasks
We begin by showing how state-of-the-art language models fail to produce diffuse distributions on simple tasks, even when a user’s prompt requests randomness. We consider two tasks: random baby name generation (BabyNames) and random number generation (RNG). In BabyNames, the model is instructed to generate a random baby name in English. In RNG, we ask the model to generate a random number between 1 and 10. The tasks have the following prompts:
@@
We consider three state-of-the-art language models: Gemma-7B-instruct (Google, 2024), Llama-2-13B-chat (Touvron et al., 2023) and Mistral-7B-instruct (Jiang et al., 2023). We use chat/instruct versions of the models, which are fine-tuned on dialog data to make them more suitable than base models for our test cases, which require instruction following. In all experiments, we sample from models using a temperature of one, which is equivalent to sampling from the maximally diffuse, unmodified model distributions. [^6] We could technically set temperature to beyond one to smooth out the model distribution further, but this is rarely done in practice due to degeneration. We also note that the untuned, baseline models need to be sampled at a temperature $>10$ to match the entropy of our fine-tuned models in Section 4.2. For each task, we sample 1,000 generations from the model and report coverage and entropy statistics of the empirical distributions of model generations.
Task Model Entropy $\uparrow$ Coverage-1000 $\uparrow$
Baseline Tuned Natural Baseline Tuned Natural
Gemma 1.16 2.28 8 10
RNG Llama-2 1.07 2.29 2.30 7 10 10
Mistral 1.88 2.30 9 10
Gemma 3.19 5.69 170 442
BabyNames Llama-2 3.24 5.83 6.43 129 490 714
Mistral 4.02 5.78 225 479
@@ 4.1 Language models fail to follow instructions for randomness
Figure 1 shows output distributions of several untuned, baseline models for BabyNames and RNG. Despite the prompts asking for randomness, these output distributions are extremely imbalanced.
For RNG, it is easy to see that the empirical distribution is far from the uniform distribution we would expect. For BabyNames, it is less clear what we should expect from the model. Do we want the generated names to be a uniform sample from the set of all names? Or do we want the names to reflect the distribution of baby names among recently born babies in the Unites States? To address this, we introduce the concept of a natural distribution for each task. For RNG, this is simply the uniform distribution; for BabyNames, we compare to a sample we draw from data on names used for babies born in the United States in the year 2022, using birth data from the US Social Security Administration. [^7] https://www.ssa.gov/OACT/babynames We note, that we are only using the natural distributions as reasonable baselines for comparison.
@@ 4.2 Optimizing for BabyNames and RNG diversifies generation
In these experiments, we see our methods can very effectively align the model distribution with the ideal distributions (Table 1). With RNG, tuned models produce near-uniform distributions over the valid choices with full coverage and near perfect entropy. With BabyNames, where the training set covers only a small portion of the possible outputs, we see an encouraging indication of generalization: while the coverage remains smaller than that of the natural distribution, it is over twice the baseline for all models and the majority of outputs are not among targets in the fine-tuning set.
@@
While RNG and BabyNames serve as toy problems, these settings have a major limitation—the fine-tuned models are evaluated on the same prompts and sample spaces in training. To further test the generalization performance of our models, we also consider settings where the prompt and sample space differ between training and evaluation. We fine-tune Llama-2 to produce diffuse probabilities over two ranges of random integers—one to ten and one to 100—and test its generalization to different instruction formats and unseen sample spaces (see Appendix B.1 for prompts). In Figure 2(a), we observe encouraging generalization trends: the tuned model produces near uniform distributions for unseen prompt formats over number ranges not in the fine-tuning set, for example from 1 to 45. When we vary the size of the random sample space of RNG (e.g., 154 to 204), the tuned model still produces substantially higher entropy distributions than the untuned model (Figure 2(b)). With these first experiments, we demonstrate that tuning models for diffuse probabilities is a promising method for increasing generation diversity.
@@ 5 Generalization Across Tasks
We use leave-one-out experiments on a set of six tasks to show that models tuned for diffuse probabilities do, in fact, have strong transferability to tasks unseen during tuning. We consider the following six tasks. Each task has several associated prompts, one of which is shown below:
@@
In Figure 3, we report in-distribution (ID) and out-of-distribution (OOD) results, which correspond to tuning sets that include or exclude the particular task, respectively. The results show a convincing trend: for all three models, our tuning method led to substantial improvements in entropy over the baselines, even when the task was held out from the tuning set. [^8] Coverage results show similar trends, and we report them in Table 5, Appendix C.1. Another interesting observation is that the baseline Mistral model consistently produces more diffuse distributions than the baseline Gemma and Llama-2 models, but after tuning, all three models have comparable entropy.
In two of the tasks (Countries and Fruits), we observe sizable generalization gaps between in-distribution and out-of-distribution entropy, which suggest that task-specific tuning remains useful, especially when we can come up with a reasonbly diverse set of generation targets for fine-tuning. However, coming up with a large enough target set isn’t always easy. In these cases, we rely on the generalization of the models trained on a diverse set of tasks. For example, in Occupations, the authors could only come up with a small set of 17 professions that start with the letter “A,” and Llama-2 and Mistral (not trained on Occupations) generalize out-of-distribution to occupations beyond what we provide in the fine-tuning set. [^9] For example, Llama-2 generated “Aromatherapist”, and Mistral generated “Agronomist.” Neither is among the fine-tuning targets of Occupations. Notably, our fine-tuning method does not substantially change the general capabilities (e.g., writing and reasoning) of the models demonstrated by evaluations on MT-Bench (see Appendix C.4), making our method compatible with tuning general-purpose large language models.
(a) Coverage results for categorical attributes.
(b) Unigram diversity of biographies.
@@ 6 Constructing More Diverse Synthetic Datasets
6.1 Out-of-distribution training improves synthetic dataset diversity
We first consider a Llama-2-13B model fine-tuned on all tasks in Section 5 and not specifically for biographies (Tuned-OOD), and compare its generations against the baseline Llama-2 model over 1000 samples. In Figure 4, we report coverage results for categorical attributes (e.g., Birth year), and normalized unigram diversity for achievements and over the entire biography. [^10] We report prompt used for biography generation and side-by-side qualitative examples in Appendix B In Table 2, we report the most frequently generated values for categorical attributes, along with the frequencies out of 1000 generations. [^11] A table of coverage and entropy statistics of generated biographies can be found in Appendix C.2.
The results on the baseline Llama-2 model indicate that the biases we observe in Section 4.1 towards certain names and numbers expectedly show up in generation. For example, out of 1000 biographies generated by the baseline Llama-2 model, 284 have the first name ’Evelyn,’ and 966 are females. Such a high level of repetition makes the resulting dataset basically unusable without substantial human intervention. In constrast, our model significantly improves the diversity of generated biographies, despite not trained specifically for generating biographies. For example, we see an $>$2X increase in coverage for most categorical attributes. There are also substantial improvements in generation diversity for achievements and over the entire biography (Figure 4(b)), although our training does not optimize for open-ended text generation.
The top five most frequent values for attributes (Table 2) help contextualize these improvements in diversity: there are significant reductions in biases towards certain attribute values. E.g., the frequency of the name “Evelyn” decreased by 25X, and the birth year 1985 by over 3X.
@@ 6.1 Out-of-distribution training improves synthetic dataset diversity
6.2 Diversifying distributions of categorical attributes
At a comparable level of generation diversity to the Tuned-OOD model (Figure 4), the Tuned-ID model is able to generate biographies with much more balanced distributions of categorical attributes: Table 2 shows that both gender and birth years are roughly uniformly distributed, and a wider range of birth places are produced by the model. [^12] See maps of generated birth places for all three models in Figure 5, Appendix C. Crucially, the model remains highly diverse on open-ended generation of achievements even when being trained exclusively on categorical attributes. This result highlights the potential in extending our fine-tuning method towards improving diversity in open-ended text generation.